OpenStudy (keana):

What are the approximate solutions of 4x^2 + 3 = -12x to the nearest hundredth? a:x ≈ -3.23 and x ≈ 0.23 b:x ≈ -2.72 and x ≈ -0.28 c:x ≈ 0.28 and x ≈ 2.72 d:x ≈ -0.23 and x ≈ 3.23

OpenStudy (anonymous):

They probably want you to use the quadratic formula.

OpenStudy (anonymous):

The quadratic formula is long and seems complicated, but it's really not so bad! Simply write your equation in this form Ax^2 + Bx + C = 0 and then plug A,B,and C into the forumla. Easie peasie parchesie. =D

OpenStudy (anonymous):

first, move that -12x to the other side and set it all to 0 4x^2 + 12x + 3 = 0 now factor (through w/e factoring method you use) and that will give you your answer :P

OpenStudy (keana):

is it x=-0.2753, -2.7247 ?

OpenStudy (anonymous):

Yes. Good job =)

OpenStudy (anonymous):

probably, then "B" would be your answer :P as long as you plugged everything in right and followed it correctly, it should be right

OpenStudy (anonymous):

Also, I wouldn't listen to harharf. Factoring here is really not possible because the roots aren't rational. Just use the quadratic formula, which it looks like you did correctly =)

OpenStudy (anonymous):

psssh, i said factor by w/e method you use (ie: decomp, educated quessing, or quadratic formula) ^_^